a cylinder has a diameter of 20cm and a lenght of 20cm.what is its volume?
Answer by checkley77(12844)

Base area=pi*r^2=3.14(20/2)^2=3.14*10^2=31.4 cm^2.
31.4*20=628 cm^2 is the volume.